HUBBLE SPACE TELESCOPE DAILY REPORT #2165 PERIOD COVERED: 0000Z (UTC) 07/14/98 - 0000Z (UTC) 07/15/98 Daily Status Report as of 196/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in yesterday's report and in HSTAR 6584, the acquisition for the first iteration of this proposal failed when the search radius limit was exceeded on FGS-2. This immediately followed the full maneuver update failures described in HSTAR 6583. One observation was affected when the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As described in yesterday's report and in HSTAR 6584, the acquisition for the first iteration of this proposal failed when the search radius limit was exceeded on FGS-2. This immediately followed the full maneuver update failures described in HSTAR 6583. One observation was affected when the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed Two Sets of NIC/1/2/3 7820 (The Metallicity Distribution of the Globular Cluster Systems of Giant Elliptical Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to study the metallicity distribution in the globular clusters contained in the large elliptical galaxies NGC-4073 and IC4951. This is an ideal use of infrared data to significantly augment existing optical data. The observations were completed as scheduled, and no anomalies were reported. 1.4 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.5 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroid HER1-13088. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD/MA1/MA2 7269 (The 1997/98 Eclipse of VV Cephei) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make ultraviolet spectrographic observations of the long-period eclipsing binary star HD208816 (a.k.a., VV-Cephei). This eclipse will provide an unprecedented opportunity to obtain spatially resolved information about the density, ionization and velocity structure in the outer atmosphere of the M supergiant primary. A wave length calibration observation was also taken. The observations were completed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.9 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/1/2/3 7225 (Proper Motion of Stars Near SGR-A) The Near Infrared Camera (Camera One, Two and Three) was used to make astrometric proper motion observations of stars in the vicinity of Sagittarius-A (a.k.a., NGC-4147). Sgr-A may be a black hole with a mass close to a million solar masses. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy QDOTULIR16. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7869 (The Morphology of Dense Gas in Seyferts , Obstruction and Fueling of AGNs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Seyfert galaxy Markarian1 to study the morphology of its dense gas. Near-Infrared color maps and images in H-2 emission will be used to trace hot and cold dense gas in this Seyfert galaxy. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Operations Notes: The SSR EDAC error counter was cleared three time using ROP SR-1. A real-time map of FHST trackers 1 and 2 was conducted at 195/1458Z. As directed by Operations Request 13878, the Battery 1 Capacity Check Test began with battery discharge at 195/2231Z. Discharge will terminate nominally at 15 volts by the Special Hardware Limit Test, expected to occur approximately 16 to 18 hours after the onset of discharge. The DF-224 SPC memory was successfully dumped to CCS string A at 196/0508Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH